OpenStudy (anonymous):

Simplify the square root of 32/5

OpenStudy (anonymous):



OpenStudy (venomblast):

you can split the the square root of 32 into \[\sqrt{16} and \sqrt{2}\]. square root of 16 is 4. so you will have \[4\sqrt{2}/\sqrt{5}. now we have a radical on the denominator. we dont want that. you will need to rationalize.

OpenStudy (venomblast):